Assume Dick's Sporting Goods sells three styles of youth hockey sticks:



Multiple Product Break-Even Analysis

Assume Dick's Sporting Goods sells three styles of youth hockey sticks: Bauer, Warrior and CCM. Presented is information for Dick's three products. 


BauerWarriorCCM
Unit selling price$180$120$100
Unit variable costs1207560
Unit contribution margin$60$45$40

With monthly fixed costs of $150,000, the company sells two Bauer sticks for each Warrior, and three Warrior for each CCM. 

Determine the number of Warrior sticks sold at the monthly break-even point. 

Round UP answer to the nearest unit. 
Answer #1

Calculation of sales mix

2 Bauer = 1 Warrior

3 Warrior = 1 CCM

Sales Mix: Bauer: Warrior: CCM = 6: 3: 1

Sales Mix in % = 60%: 30%: 10%

Contribution Mix = $60: $45: $40

Weighted average contribution Mix = 60*60% + 45*30% + 40*10%

=$36 + $13.5 + $4 = $53.5

Fixed cost (Given) = $150,000

Break Even Point (Units) = Fixed Cost / Weighted average contribution Mix

=$150,000 / $53.5 = 2804 Units

No of Warriors sticks sold at the monthly break even points

=2804*30% = 841 units

Please like this answer it helps me a lot. Thank You.

> This was very helpful!! thank you so much!

Karen Cerrato Hernandez Wed, Nov 17, 2021 1:09 PM